OpenStudy (anonymous):

Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1

OpenStudy (zzr0ck3r):

\(f(g(-1)) =f((-1)+3)=f(2) = 5(2)-10 = \ \large ? \)